Find the class boundaries, midpoints and widths for each class.



1. 58-62


2.125-131


3.16.35-18.46


4.16.3-18.5

1
Expert's answer
2021-12-21T02:35:53-0500

1. 58-62

Lower class boundary = 58, Upper class boundary = 62

Mid point = (58+62)/2 = 60

Width of the class = 62 - 58 = 4


2.125-131

Lower class boundary = 125, Upper class boundary = 131

Mid point = (125+131)/2 = 128

Width of the class = 131 - 125 = 6


3.16.35-18.46

Lower class boundary = 16.35, Upper class boundary = 18.46

Mid point = (16.35+18.46)/2 = 17.405

Width of the class = 18.46 - 16.35 = 2.11


4.16.3-18.5

Lower class boundary = 16.3, Upper class boundary = 18.5

Mid point = (16.3+18.5)/2 = 17.4

Width of the class = 18.5 - 16.3 = 2.2